OpenStudy (anonymous):

In a group of 760 people, 510 can speak Hindi and 360 can speak English. Find (i) How many can speak both Hindi and English? (ii) How many can speak Hindi only? (iii) How many can speak English? i need steps how to solve this, not just answers.

OpenStudy (anonymous):

let say hindi H and english E H=510 E=360 510+360=870 but we have 760 people and 870 - 760=110 this mean 110 people know both of these.. 510-110=400 this is numbers of people knows only hindi 360-110=250 knows only english

OpenStudy (anonymous):

is that enough?

OpenStudy (anonymous):

yah thank u soooo much:)

OpenStudy (anonymous):

and how about this one? On a rainy day, 4/15 of the strength of a class was absent and 1/2 of those present, brought raincoat.How many did not bring raincoat, the total strength of the class being 90

OpenStudy (anonymous):

as far as i understand, 4/15 absent and 11/15 present brought raincoats number is 1/2 of 11/15 , it is 11/30 of class then 4/15=8/30 absent and 11/30 brought raincoat multply these 8/30+11/30=19/30 90*(19/30)=57 90-57=33 this is didnt brings raincoat's number ;)

OpenStudy (anonymous):

:) thank you :) more plzz...:$ A man bought banana at 5 for RS 4 and sold at 3 for RS 2. Find his gain or loss percent.

OpenStudy (anonymous):

and the last one plz A fruit seller bought 240 oranges at RS 20 per dozen. He sold the same at RS 2 per orange.Find the gain or loss percent